# Diversified and Adaptive Negative Sampling on Knowledge Graphs

### v1

@@ 4.3 Additional research questions
(a) Ablation study on generator and FiLM
@@
Parameter sensitivity (RQ3). To answer RQ3, we perform a parameter sensitivity analysis. We first analyze how the number of negative triplets per positive triplet, $N_{s}$, can impact model performance. As shown in Figure 2(b), as we increase $N_{s}$ on each dataset, we consistently observe that the MRR performance improves and peaks at $N_{s}=20$. A larger $N_{s}$ allows for greater diversity, which explains the initial increase in performance. However, when $N_{s}\geq 20$, performance starts to plateau or even deteriorate, due to highly imbalanced training data.
@@
Case study (RQ4). We conduct a qualitative evaluation of DANS to demonstrate the diversity and adaptiveness of negative triplets generated by DANS.
@@
Figure 3: Visualization of diversity comparing to Random Negative Sampling (RNS). Best viewed in color.
(a) $\mathtt{hasPart}$
(b) $\mathtt{animalType}$
(NELL-995)
(c) $\mathtt{isA}$
Figure 4: Visualization of diversity comparing to Popularity-weighted Negative Sampling (PNS). Best viewed in color.
Note that the ablation study has demonstrated improved model performance with the two-way generator, FiLM layers and provided direct evidence on the importance of diverse pathways and FiLM in the model architecture. However, it is not immediately clear if having two pathways and FiLM layers in the generator would indeed produce diverse and adaptive examples in their embedding space. Instead, as asked in RQ4, can we observe some generated examples on their diversity and adaptiveness?
@@
To demonstrate the notion of diversity in RQ4, we present a few case studies of how DANS can produce more diverse examples than uniform random sampling (RNS) or popularity-weight random sampling (PNS). In Figures 4 and 4 we visualize the positive and negative tail entities w.r.t. a given relation and all its head entities on each dataset. More specifically, each point represents one tail entity, which can be a positive (real) tail entity, or a negative tail entity. The negative entity can be generated by one of the pathways $G_{E}$ or $G_{R}$ of the generator, or randomly sampled by RNS or PNS. The high-dimensional embedding space is projected onto a Cartesian plane using the $t$-SNE algorithm [37]. In Figure 4, we compare the diversity of negative entities generated by DANS with that of RNS-based negative entities. For our case study, we select one relation for each dataset, namely, $\mathtt{hasPart}$ on WN18RR, $\mathtt{animalType}$ on NELL-995 and $\mathtt{isA}$ on UMLS, so that all the positive and negative tail entities for a common relation (and the same original head entities) can be contrasted in one visualization. The results show that DANS could provide more diverse negative entities for model training, where those generated by $G_{E}$ and $G_{R}$ occupy different subspaces from the positive entities. In contrast, RNS lacks diversity and samples negative entities in the same subspace as positive entities. This could even potentially contribute to false negative triplets as they are not well separated from the real ones. Similarly, in Figure 4, we compare the negative entities generated from $G_{E}$ and $G_{R}$ in DANS against the output of PNS, which replaces uniform negative sampling with popularity-weighted sampling. The results again echoed that DANS has produced more diverse negative entities in different subspaces as the positive entities, whereas PNS samples negative entities that mostly overlap with the positive triplets with less diversity.
(a) Generated samples w/o FiLM
(b) Generated samples w/ FiLM
Figure 5: Visualization of adaptiveness for relations $\mathtt{DomainRegion}$ and $\mathtt{DomainUsage}$ in WN18RR. (a) Generated samples w/o FiLM layers and (b) w/ FiLM layers. Best viewed in color.
On the other hand, to demonstrate the notion of adaptiveness, we compare two different relations from each dataset to visualize its differences with and without FiLM layers. For example, Figure 5 visualizes the impact of adaptiveness for relations $\mathtt{DomainRegion}$ and $\mathtt{DomainUsage}$ in the WN18RR dataset. “Positive1” and “Positive2” denote the existing positive entities of each of the two relations, respectively; “Fake1” and “Fake2” denote the corresponding negative samples produced from the generators for each of the two relations, respsectively. As shown in Figure 5(a), without FiLM layers, “Fake1” and “Fake2” both spread across the plane without any clear association to the corresponding “Positive1” and “Positive2”, and thus offering less discriminative power to improve the learning of each relation. In contrast, in Figure 5(b), after FiLM layers are added, we can clearly visualize that “Fake1” and “Fake2” are adapted to “Positive1” and “Positive2” as they move closer to each of their respective positive samples, improving the discriminative power of learning. This demonstrates that the global generators are adapted into local models conditioned on individual input (entity or relation) when FiLM layers are present. Similar patterns can be observed in NELL-995 and UMLS datasets, which are presented in E.
5 Conclusion and Future Work
@@
In this work, we introduced DANS, a negative sampling strategy for knowledge graph embedding that explicitly accounts for the informativeness of negative triplets. On one hand, we proposed a two-way generator to increase the overall informativeness by diversifying the negative triplets based on their association with not only entities but also relations. On the other hand, we adapt the global generator model into local models, which generate negative triplets in a finer-grained manner to improve their individual informativeness. Empirically, DANS has outperformed state-of-the-art baselines on three benchmark knowledge graphs through both quantitative and qualitative experiments. In the future, we believe that the concept of diversity and adaptiveness can be further extended to other graph representation learning problems with a lean structure that is less parameter intensive.
6 Funding
This research / project is supported by the Ministry of Education, Singapore, under its Academic Research Fund Tier 2 (Proposal ID: T2EP20122-0041). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of the Ministry of Education, Singapore.
Appendices
Appendix A Scoring function
ComplEx: ComplEx [36] extends DistMult [44] by introducing complex-valued embeddings to better model asymmetric relations. In ComplEx, entity and relation embeddings no longer lie in a real space but a complex space $\mathbb{C}^{k}$ by operator $\operatorname{Re}$.
$\mathcal{F}(\langle h,r,t\rangle)_{\text{ComplEx}}=\operatorname{Re}(\mathbf{e}_{h}^{\top}\text{Diag}(\mathbf{e}_{r})\mathbf{e}_{t})$
ComplEx maps the entities and relations to the complex vector where $\mathbf{e}_{h},\mathbf{e}_{t}\in\mathbb{C}^{k}$ by operator $\operatorname{Re}$ and $\text{Diag}(\mathbf{e}_{r})\in\mathbb{R}^{d\times d}$ is diagonal matrix whose diagonal is $\mathbf{e}_{r}$, an $r$-specific trainable vector of the decoder.
RotatE: Inspired by TransE [2], RotatE [34] veers into complex vector space and is motivated by Euler’s identity. The model defines each relation as a rotation and measures how the distance from the source entity to the target entity to account for three relation patterns: symmetric/anti-symmetric, inversion and composition.
$\mathcal{F}(\langle h,r,t\rangle)_{\text{RotatE}}=-||\mathbf{e}_{h}\ \mathcal{\circ}\ \mathbf{e}_{r}-\mathbf{e}_{t}||^{2}$
RotatE maps the entities and relations to the complex vector where $\mathbf{e}_{h},\mathbf{e}_{t}\in\mathbb{C}^{k}$ and $\circ$ denotes the Hadamard product.
Appendix B Evaluation Metrics
Mean Reciprocal Ranking (MRR): For each candidate list we record the reciprocal of the ranking position of the positive triplet, and take the average over all lists. It reflects the absolute ranking of the first relevant item in the list.
$\operatorname{MRR}=\textstyle\frac{1}{|Q|}\sum_{i=1}^{|Q|}\frac{1}{\operatorname{rank}_{i}}$
Where $rank_{i}$ is the position of the relevant result in the ith query and Q is the total number of queries.
Hit ratio at $K$ (H@$K$): We compute the fraction of candidate lists in which positive triplets fails within the first $K$ positions.
$\operatorname{HR}=\textstyle\frac{\left|U_{\text{hit }}^{K}\right|}{\left|U_{\text{all }}\right|}$
where $\left|U_{\text{hit }}^{K}\right|$ is the number of times a positive triplet is ranked within top $K$ positions in the candidate list. $\left|U_{\text{all }}\right|$ is the total number of candidates.
Normalized discounted cumulative gain at $K$ (NDCG@$K$): We compare the ranked list with the ideal list, where a match at a lower position would have a discounted gain. The gain is further normalized across lists, and we measure the average over all lists. It reflects the relevance at the top $K$ positions, taking position significance into account.
$\mathrm{DCG}_{K}=\textstyle\sum_{i=1}^{K}\frac{Gain}{\log_{2}(i+1)}$
$\mathrm{IDCG}_{K}=\textstyle\sum_{i=1}^{K}\frac{Gain}{\log_{2}(i+1)}$
$\mathrm{nDCG}_{K}=\textstyle\frac{\mathrm{DCG}_{K}}{\mathrm{IDCG}_{K}}$
where $\mathrm{DCG}_{K}$ measures our ranked list and $\mathrm{IDCG}_{K}$ measures the ideal list. The numerator $Gain$ equals to 1 when prediction of positive triplet falls within the top $K$ positions and 0 if otherwise. In addition, $i$ refers to the ranking position of the positive triplet.
Appendix C Efficiency Comparison
Table 4: Training time comparisons (minutes) for different negative sampling strategies with the RGCN backbone and DistMult decoder on the WN18RR dataset.
{tabu}
l *4c Number of Negative Samples
Model 50 20 105
RGCN 184 87 79 68
RGCN-P 182 94 79 67
RGCN-Adv 191 108 91 70
RGCN-MCNS 325 282 169 137
DANS 761 461 271 201
To compare the efficiency of different negative sampling methods, we report the runtime of DANS and other negative sampling strategies including random negative sampling in RGCN, popularity biased negative sampling in RGCN-P, self-adversarial negative sampling in RGCN-Adv and negative sampling via positive distribution in RGCN-MCNS in Table C. We standardize the experimental setting and use the same RGCN backbone with DistMult scoring function for fair comparison. While DANS incurs more time than other negative sampling methods, the increase over more advanced methods like RGCN-MCNS is by a manageable constant factor. Furthermore, the growth in time is linear when more negative samples are generated.
Using 20 negative samples for reference, results show that RGCN-Adv, that weighs the negative triplets takes 108 minutes to complete, which is slightly more than 87 minutes and 94 minutes in RGCN and RGCN-P. RGCN-MCNS which samples from positive distribution requires more time (135 minutes) to finish. Overall, DANS being a more complexed model that generally performs better than baselines, takes approximately 3 to 5 times longer to run. As the number of negative samples increases, the gap for run-time widens between DANS and other model variants as DANS takes more computational resources to generate negative synthetic samples.
Appendix D Performance on the FB15k-237 dataset
Table 5: Performance comparison with baselines (all using the ComplEx decoder). See Table II caption for entry
FB15k-237
Model MRR H@1 H@5 H@10 NDCG@5
ComplEx .211$\pm$.005 .128$\pm$.003 .234$\pm$.006 .381$\pm$.006 .128$\pm$.004
SANS-RW .195$\pm$.006 .153$\pm$.004 .204$\pm$.08 .267$\pm$.008 .119$\pm$.004
NSCaching .261$\pm$.007 .183$\pm$.006 .284$\pm$.007 .417$\pm$.006 .192$\pm$.008
KBGAN .259$\pm$.008 .169$\pm$.0106 .279$\pm$.007 .413$\pm$.010 .188$\pm$.006
CAKE .195$\pm$.007 .132$\pm$.004 .222$\pm$.006 .319$\pm$.009 .130$\pm$.003
SMiLE .179$\pm$.003 .114$\pm$.003 .197$\pm$.005 .284$\pm$.007 .112$\pm$.004
DANS .253$\pm$.007 .185$\pm$.006 .333$\pm$.009 .446$\pm$.011 .252$\pm$.007
The DistMult decoder we use on the other datasets tends to perform poorly on FB15k-237 in Dettmers et al. [7], Ji et al. [14], Zhou et al. [51] Hence, on FB15k-237, we implement the baselines and DANS using the ComplEx decoder, as shown in Table 5. Results show that DANS still achieves competitive performance in comparison to the baselines. In particular, DANS achieves a notable gain of 17.3%, 6.95% and 31.3% in terms of H@5, H@10 and NDCG@5, respectively.
Appendix E Case Study for Adaptiveness
(a) Generated samples without FiLM
(b) Generated samples with FiLM
Figure 6: Visualization of adaptiveness for relations $\mathtt{Animal.eat.vegetable}$ and $\mathtt{sports.school.in.country}$ on NELL-995. Best viewed in color.
(a) Generated samples without FiLM
(b) Generated samples with FiLM
Figure 7: Visualization of adaptiveness for relations $\mathtt{Domain\_Location\_of}$ and $\mathtt{Domain\_Property\_of}$ on UMLS. Best viewed in color.
We conduct additional experiments to qualitatively assess the impact of FiLM layer on output from the generator pathways through a case study. To demonstrate adaptiveness, we compare two different relations from each dataset to visualize its differences with and without FiLM layers. The visualization on the WN18RR dataset has already been explained in the main paper. Here, we include the visualizations on the NELL-995 and UMLS datasets in Figures 7 and 7, respectively, which display similar patterns as explained in the main paper.

### v2

@@ 4.3 Additional research questions
(a) Ablation study
@@
Parameter sensitivity (RQ3). To answer RQ3, we perform a parameter sensitivity analysis. We first analyze how the number of negative triplets per positive triplet, $N_{s}$, can impact model performance. As shown in Figure 2(b), as we increase $N_{s}$ on each dataset, we consistently observe that the MRR performance improves and peaks at $N_{s}=20$. A larger $N_{s}$ allows for greater diversity, which explains the initial increase in performance. However, when $N_{s}\geq 20$, performance starts to plateau or even deteriorate, due to imbalanced training data.
@@
Figure 3: Visualization of diversity. Best viewed in color.
@@
Case studies (RQ4). We conduct a qualitative evaluation of DANS to demonstrate the diversity of negative triplets generated by DANS. While the ablation study has demonstrated the importance of diversity through improved model performance with the two-way generator, we further present a few case studies of how DANS can produce more diverse examples than uniform random sampling (RNS). In Figures 3, we visualize the positive and negative tail entities w.r.t. a given relation and all its head entities on each dataset. More specifically, each point represents one tail entity, which can be a positive (real) tail entity, or a negative tail entity. The negative entity can be generated by one of the pathways $G_{E}$ or $G_{R}$ of the generator, or randomly sampled by RNS. The high-dimensional embedding space is projected onto a Cartesian plane using the $t$-SNE algorithm [16]. In Figure 3, we compare the diversity of negative entities generated by DANS with that of RNS-based negative entities. For our case study, we select one relation for each dataset, namely, $\mathtt{hasPart}$ on WN18RR, $\mathtt{animalType}$ on NELL-995 and $\mathtt{isA}$ on UMLS, so that all the positive and negative tail entities for a common relation (and the same original head entities) can be contrasted in one visualization. The results show that DANS could provide more diverse negative entities for model training, where those generated by $G_{E}$ and $G_{R}$ occupy different subspaces from the positive entities. In contrast, RNS lacks diversity and samples negative entities in the same subspace as positive entities. This could even potentially contribute to false negative triplets as they are not well separated from the real ones.
5 Conclusion
@@
In this work, we introduced DANS, a negative sampling strategy for knowledge graph embedding that explicitly accounts for the informativeness of negative triplets. On one hand, we proposed a two-way generator to increase the overall informativeness by diversifying the negative triplets based on their association with not only entities but also relations. On the other hand, we adapt the global generator model into local models, which generate negative triplets in a finer-grained manner to improve their individual informativeness. Empirically, DANS has outperformed state-of-the-art baselines on three benchmark knowledge graphs through both quantitative and qualitative experiments.